# Extending Congruences for the number of smallest parts in overpartitions with smallest part even

### v1

@@ 1. Introduction
We recall that an overpartition [3] is a partition in which the first occurrence of each part may be overlined. Garvan and Jennings-Shaffer [5] introduced the function $\overline{spt2}(n)$, which is the number of smallest parts in the overpartitions of $n$ where the smallest part is even and is not overlined. For instance, considering the 14 overpartitions of 44, namely
@@
Our goal in this paper is to extend the congruences modulo 44 satisfied by $\overline{spt2}(n)$. This is done in two directions. First, we prove some new individual congruences for $\overline{spt2}(n)$. Then, we provide a number of infinite families of Ramanujan-like congruences satisfied by $\overline{spt2}(n)$. The new individual congruences are the following.
Theorem $1.$
For all $n\geq 0$.
@@
(4) $\overline{spt2}(64n+56)\equiv 0\pmod{4}$,
(5) $\overline{spt2}(72n+42)\equiv 0\pmod{4}$,
(6) $\overline{spt2}(80n+34)\equiv 0\pmod{4}$,
(7) $\overline{spt2}(80n+66)\equiv 0\pmod{4}$.
@@
(11) $\overline{spt2}(2^{j+3}9n+21\cdot 2^{j+1})$ $\equiv 0\pmod{4}$
@@
This paper is organized as follows. In Section 2, we recall some basic identities that are used in the proof of Theorem $1.$ Section 3 is devoted to the proof of Theorem $1.$ The proofs of the infinite families of Ramanujan–like congruences shown in Theorem 1.2 are presented in Section 4. The proof of Theorem 1.3 is the content of Section 5.
@@ 2. Preliminaries
(15) $\psi(q)$ $:=f(q,q^{3})=\sum_{n=0}^{\infty}q^{n(n+1)/2}=\frac{(q^{2};q^{2})_{\infty}^{2}}{(q;q)_{\infty}}.$
These functions satisfy many interesting properties (see Entries 18, 19, and 22 in [1]), including:
@@
In order to prove (4), we recall the following (see [8, Eq. (4.16)]):
Using (18), we get
which, after dividing by $q$ and replacing $q^{2}$ by $q$, yields
Since $f_{2}^{3}f_{4}^{3}$ is a function of $q^{2}$, then (4) follows.
@@
Finally, we prove (6) and (7). In order to do so, we substitute (15) and (25) into (28) to obtain
$\sum_{n=0}^{\infty}\overline{spt2}(16n+2)q^{n}$ $\equiv 2\sum_{k=0}^{\infty}q^{k(k+1)/2}-\sum_{l=0}^{\infty}(-1)^{l}q^{l(l+1)/2}\pmod{4}$.
Since $j(j+1)/2\not\equiv 2\mbox{ or }4\pmod{5}$, the congruences (6) and (7) follow.
@@ 4.2. Proof (9)
This is Equation (2) in Theorem $1.$
@@ 4.3. Proof (10)
This is Equation (3) in Theorem $1.$
@@
This is Equation (5) in Theorem $1.$
@@ 4.5. Proof (12)
This is Equation (6) in Theorem $1.$
@@ 4.6. Proof (13)
This is Equation (7) in Theorem $1.$
@@
Using Lemma 25, it follows from (29) that
@@
Note that $n=k(k+1)+2l(l+1)$ is equivalent to writing $4n+3=(2k+1)^{2}+2(2l+1)^{2}$. Thus, $\overline{spt2}(32n+24)\equiv 0\pmod{4}$ if $4n+3$ is not of the form $x^{2}+2y^{2}$. Since $p\equiv 5\mbox{ or }7\pmod{8}$ we know that $\left(\frac{-2}{p}\right)=-1$, i.e., 2-2 is a quadratic nonresidue modulo pp. This implies that $\nu_{p}(N)$ is even if NN is of the form $2x^{2}+y^{2}$. However, here $n=p^{2k+1}m+3(p^{2k+2}-1)/{4}$. Then

### v2

@@ 1. Introduction
We recall that an overpartition [3] is a partition in which the first occurrence of each part may be overlined. Garvan and Jennings-Shaffer [5] introduced the function $\overline{spt2}(n)$, which is the number of smallest parts in the overpartitions of $n$ where the smallest part is even and is not overlined. For instance, considering the 14 overpartitions of $4$, namely
@@
Our goal in this paper is to extend the congruences modulo $4$ satisfied by $\overline{spt2}(n)$. This is done in two directions. First, we prove some new individual congruences for $\overline{spt2}(n)$. Then, we provide a number of infinite families of Ramanujan-like congruences satisfied by $\overline{spt2}(n)$. The new individual congruences are the following.
Theorem 1.1.
For all $n\geq 0$,
@@
(4) $\overline{spt2}(72n+42)\equiv 0\pmod{4}$.
@@
(8) $\overline{spt2}(2^{j+3}9n+21\cdot 2^{j+1})$ $\equiv 0\pmod{4}$,
@@
This paper is organized as follows. In Section 2, we recall some basic identities that are used in the proof of Theorem 1.1. Section 3 is devoted to the proof of Theorem 1.1. The proofs of the infinite families of Ramanujan–like congruences shown in Theorem 1.2 are presented in Section 4. The proof of Theorem 1.3 is the content of Section 5.
@@ 2. Preliminaries
(12) $\psi(q)$ $:=f(q,q^{3})=\sum_{n=0}^{\infty}q^{n(n+1)/2}=\frac{(q^{2};q^{2})_{\infty}^{2}}{(q;q)_{\infty}},$
where $(a;q)_{\infty}:=(1-a)(1-aq)(1-aq^{2})\cdots$ is the usual $q$-Pochhammer symbol. These functions satisfy many interesting properties (see Entries 18, 19, and 22 in [1]), including:
@@ 4.2. Proof of (6)
This is Equation (2) in Theorem 1.1.
@@ 4.3. Proof of (7)
This is Equation (3) in Theorem 1.1.
@@ 4.4. Proof of (8)
This is Equation (4) in Theorem 1.1.
@@ 4.5. Proof of (9)
This follows directly from [8, Theorem 1.4]. Indeed, since $k(k+1)/2\not\equiv 2\pmod{5}$, we see that
$\overline{spt2}(80n+34)=\overline{spt2}(16\cdot(5n+2)+2)\equiv 0\pmod{4}$
by [8, Eq. (1.4)].
@@ 4.6. Proof of (10)
This follows directly from [8, Theorem 1.4]. Indeed, since $k(k+1)/2\not\equiv 4\pmod{5}$, we see that
$\overline{spt2}(80n+66)=\overline{spt2}(16\cdot(5n+4)+2)\equiv 0\pmod{4}$,
by [8, Eq. (1.4)].
@@
Initially we recall the following congruence (see [8, Eq. (4.16)]):
Using (15), we get
which, after dividing by $q$ and replacing $q^{2}$ by $q$, yields
Using Lemma 2.6, it follows from (26) that
@@
Note that $n=k(k+1)+2l(l+1)$ is equivalent to writing $4n+3=(2k+1)^{2}+2(2l+1)^{2}$. Thus, $\overline{spt2}(32n+24)\equiv 0\pmod{4}$ if $4n+3$ is not of the form $x^{2}+2y^{2}$. Since $p\equiv 5\mbox{ or }7\pmod{8}$ we know that $\left(\frac{-2}{p}\right)=-1$, i.e., 2-2 is a quadratic nonresidue modulo $p$. This implies that $\nu_{p}(N)$ is even if $N$ is of the form $2x^{2}+y^{2}$. However, here $n=p^{2k+1}m+3(p^{2k+2}-1)/{4}$. Then
@@
6. Concluding remarks
In the work above, we have accomplished our primary goal, which was to prove new Ramanujan–like congruences satisfied by $\overline{spt2}(n)$ modulo $4$, including some infinite families of congruences. With that said, we admit that our results above are not exhaustive; computational evidence indicates that there are more Ramanujan–like congruences modulo $4$ satisfied by $\overline{spt2}(n)$ which are not covered by our results above. We leave it to the interested reader to consider proving additional Ramanujan–like congruences which are satisfied by $\overline{spt2}(n)$.